Circular No. 6674

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

              URL http://cfa-www.harvard.edu/cfa/ps/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1997cq IN ANONYMOUS GALAXY The McDonald Observatory Supernova Search Team (M. T. Adams,

D. A. Howell, M. H. Ward, J. C. Wheeler, and W. Wren) report their discovery of an apparent supernova on three Kron-Cousins R-band CCD images of Abell 2199 (z = 0.0299) taken during June 2.274–2.286 UT with the 0.76-m telescope. SN 1997cq is located at α = 16^{h}29^{m}50^{s}.86, δ = +39^{o}45′20′′.3 (equinox 2000.0), which is 39′′ south of the nucleus of a barred spiral galaxy. Preliminary CCD photometry between June 4.146 and 4.226 yields V = 19.23, B − V = +0.13, V − R = +0.09, V − I = +0.36. SN 1997cq is not visible on R-band images (limiting mag ∼ 21) taken on 1995 July 5, 1996 Mar. 12 and 29, and Oct. 7.

GRB 970508 A. Fruchter and L. Bergeron, Space Telescope Science Institute; and E.

Pian, ITESRE, CNR, Bologna, on behalf of a large collaboration, report: “The proposed optical counterpart of GRB 970508 (IAUC 6654) was imaged with the Hubble Space Telescope, using the STIS CCD (+ clear filter) in four 1250-s exposures and the NICMOS Camera 2 (+ F160W filter) in four 500-s exposures during June 2.52–2.66 and 2.67–2.74 UT, respectively. The optical source is detected in the STIS image at magnitude R = 23.10 ± 0.15. This and previous ground-based R-band observations are consistent with a power-law fading of the source after May 10 with an index of −1.1. The NICMOS image yields magnitude H = 20.3 ± 0.3, or 6.5 ± 2.0 μJy. The primary uncertainties in both these measurements are due to our imperfect knowledge of the calibration of these two new instruments. No significant evidence of nebulosity is seen in either the STIS or NICMOS imaging. A compact galaxy coincident with the optical source brighter than R = 24.3 would, by now, have produced a noticeable flattening of the power-law falloff of the source luminosity with time. Similarly, a galaxy 1′′ in diameter that is also brighter brighter than R = 25 would be visible in the image after a point-spread function is subtracted from the position of the source. Nebulosity, such as that seen surrounding GRB 970228 (Sahu et al. 1997, Nature 387, 476), would be readily apparent. The only visible galaxies within a few arcsec of the source have integrated magnitudes fainter than R = 24.5. Therefore, if the source is at the redshift (z = 0.835) of the most distant absorption system seen in the spectrum (IAUC 6655), the host galaxy must be at least several times less luminous than the knee of the Schechter galaxy luminosity function, L_{⋆}.”

1997 June 5  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1997 CBAT  Daniel W. E. Green